write and equation in point slope form of the line through the pint (-5,2) perpendicular to the line with equation 4x+3y=7
Answer by Alan3354(69443)

write and equation in point slope form of the line through the pint (-5,2) perpendicular to the line with equation 4x+3y=7
------------------
Step 1, find the slope of the line.
Step 2, lines perpendicular have a slope, m, that's the negative inverse.
Step 3, use y-y1 = m*(x-x1) where (x1,y1) is the given point.
=====================
The practice will do you good.